[Date Prev][Date Next][Thread Prev][Thread Next][Search] [Main Index] [Thread Index] [HEASARC Mailing List Archives]

CGRO biWekly Status Report

   Compton Observatory Science Report #160, Friday June 24, 1994
     Chris Shrader, Compton Observatory Science Support Center

       Questions or comments can be sent to the Compton SSC.
          Phone: 301/286-8434
          e-mail:   NSI_DECnet: GROSSC::SHRADER 
          Internet: shrader@grossc.gsfc.nasa.gov

                         Spacecraft Status

The Observatory, scientific instruments and all spacecraft
subsystems continue to function nearly flawlessly.

                 Science Support Center Activities

Guest investigator activities at the CGRO-SSC have been on the
upswing - several new IGORE (Interactive Gro Osse Reduction
Environment) users have been utilizing SSC computing and archive
facilities. Scientists utilizing EGRET and BATSE data and analysis
software have also visited the SSC this month. Interest in the
BATSE 2B catalog continues to be significant. Efforts to install
the COMPTEL COMPASS data analysis sytem at the SSC are continuing.

Preparations for the forthcoming Cycle-4 Proposal Review Meeting
have been finalized. Additionally, the CGRO Users Committee will
meet Thursday and Friday of next week. Dr. Alan Marscher (Boston
University) is a new member of the Users Committee, and Dr. Kevin
Hurley (UC Berkley) will take over charimanship of the committee
after this meeting.

                        Instrument Reports


EGRET operations were normal during this biweekly period. Delivery
of data to the GRO SSC is on schedule. A very severe, relatively
sudden thunderstorm caused some damage to the EGRET computer
systems, but it was possible to resume data reduction after about
two days. Interaction with Guest Investigators remains at a good

During standard basic analysis of EGRET data, it was discovered
that the same source had been awarded to two guest investigators by
the Phase-3 peer-review committee. There is also a serindipitous
source involved. This matter has been brought to the attention of
the project scientist for resolution. 

A summary of the current EGRET results will be presented at the
COSPAR meeting in July, 1994. Several papers summarizing the
results of phases 1 and 2 are in preparation and should be
submitted for publication soon.

OSSE operations are normal.

In viewing period 330 (10-14 June), the Z-axis target was Cyg X-3
(PI team), and the X-axis target was GT 0236+610 (Guest
Investigators M. Coe and M. Tavani).  The Sun was not available on
the OSSE scan plane, so the slewing burst response to BATSE solar
flare triggers was disabled.  In viewing period 331.5 (14-18 June),
the Z-axis target was Cyg X-3 (PI team), and the X-axis targets
were PSR 0611+22 (Guest Investigator J. Cordes) and 4U 0115+63 in
its current outburst.  The Sun was available on the scan plane, so
the slewing response was enabled.  For the first and last 24 hours
of vp 331.0, we used two detectors to monitor the spectral state of
Cyg X-1. In viewing period 332 (18 June - 5 July), the Z-axis
target is a galactic plane scan near (l,b) = (20,0) (Guest
Investigator H. Bloemen), and the X-axis targets are Cas A (PI
team) and GT 0236+610 (Guest Investigators M. Coe and M. Tavani). 
The Sun is not available on the scan plane, so the slewing was once
again disabled.  For two 24-hour periods, we will monitor Cyg X-1
in place of Cas A.


The pulsed flux from the x-ray binary 4U 0115+634 is declining, and
is expected to drop below detection threshold in the next few days.
The transient source has been detected by BATSE for most of the
last 60 days.

The BATSE pulsed source monitor is currently (TJD 9424) detecting
pulsed flux  from 4U0115+634, Cen X-3, OAO 1657-415, Vela X-1 and
GX 301-2. This monitor is based on epoch folding of the BATSE 20-50
keV, 1.024 s resolution discriminator rates. Astronomers with
correlative observing programs that would be interested being
e-mailed the daily monitor report should contact Mark H. Finger at
Finger@BATSE.msfc.nasa.gov or BATSE::Finger.

BATSE folded-on-board data is currently being collected for the
Crab, 4u0115+634, and Her X-1.

As of June 20th, BATSE has detected 1020 cosmic gamma-ray bursts
out of a total of 2931 on-board triggers in 1154 days of operation.
There have been 733 triggers due to solar flares with emission
above 60 keV.


The COMPTEL instrument is performing well and continues routine

Several members of the team will be participating in the upcoming
CGRO Cycle 4 Proposal Review next week near GSFC.

It has been a busy (but somewhat disappointing) week of
gamma-ray-burst-related activity within the COMPTEL team.  A weak
cosmic gamma-ray burst was recorded within the field of view of
COMPTEL on Sunday, 19 June. The analysis of data from this burst
was complicated by the fact that it occurred very near the Earth's
limb, and was consequently highly contaminated by background
events. A second burst from the same general direction, reported by
BATSE on 22 June, was not seen by COMPTEL, primarily due to the
high level of background observed in this direction during the
current pointing. Finally, the COMPTEL rapid-response team was
notified automatically by the BATSE BACODINE system of a third
gamma-ray burst (in a different direction of the sky) within the
field of view of the instrument early on Thursday, 23 June;
unfortunately, no COMPTEL signal at MeV energies was measured for
this last event.

Summary information on GRB940619 follows:

NOTICE_DATE: Wed Jun 22 15:51:19 GMT 1994
ORIG_SENDER: burst@comptel.unh.edu
BURST_DATE :  9522 TJD; 19-JUN-1994
BURST_TIME : 77480.00 SEC; 21:31:20 UT
BURST_SCZEN:  27.3 Deg.
BURST_SCAZI: 110.2 Deg.
BURST_CLASS: Weak  (Approx  29 events)
COORD_J2000: [RAsc     Decl]   [RAsc            Decl]
============ ===============   ======================
MAXLIK_POS : [297.7d -31.9d] = [19h50m38s -31d55m58s]
ERRBOX1_POS: [295.1d -32.7d] = [19h40m35s -32d40m21s]
ERRBOX2_POS: [297.6d -30.3d] = [19h50m16s -30d16m41s]
ERRBOX3_POS: [299.7d -31.8d] = [19h58m40s -31d45m54s]
ERRBOX4_POS: [297.3d -34.2d] = [19h49m 3s -34d12m 2s]
COORD_B1950: [RAsc     Decl]   [RAsc            Decl]
============ ===============   ======================
MAXLIK_POS : [296.9d -32.1d] = [19h47m34s -32d 3m58s]
ERRBOX1_POS: [294.4d -32.8d] = [19h37m29s -32d47m40s]
ERRBOX2_POS: [296.8d -30.4d] = [19h47m14s -30d24m39s]
ERRBOX3_POS: [298.9d -31.9d] = [19h55m37s -31d54m27s]
ERRBOX4_POS: [296.5d -34.3d] = [19h45m55s -34d19m55s]